OpenStudy (anonymous):

Solve: cot x / tan x = csc^2 x -1 completely lost on how to solve or where to start

OpenStudy (anonymous):

looking up ratios now

OpenStudy (mimi_x3):

\[\frac{cotx}{tanx} => \frac{cosx}{sinx} * \frac{1}{tanx} => \frac{cosx}{sinx}* \frac{cosx}{sinx} => \frac{cos^2x}{sin^2x} = \frac{1}{sin^2x} -1 \] \[ \frac{cos^2x}{sin^2x} = \frac{1-sin^2x}{sin^2x}\] \[\frac{cos^2x}{sin^2x} - \frac{1-sin^2{x}}{sin^2x} = 0 => \] \[\frac{cos^2x -(1-sin^2{x})}{sin^2x}=0\]

OpenStudy (anonymous):

wouldn't I divide the sinx at the end by it self or wont it cancel?

OpenStudy (mimi_x3):

what do you mean?

OpenStudy (anonymous):

in the last part 1 - sinx^ 2/ sinx^2 wont these two cancel out because they are the same

OpenStudy (mimi_x3):

I think I made a mistake somewhere; I should do it on paper lol \[\frac{cos^2(x)}{sin^2(x)} = \frac{cos^2x}{sin^2x}\]

OpenStudy (mimi_x3):

It's quite odd though. \[\frac{cos^2(x)}{sin^2(x)} - \frac{cos^2(x)}{sin^2(x)} =0\]

OpenStudy (anonymous):

yeah I agree

OpenStudy (mimi_x3):

Hmm..i think there is something wrong with the question; we are solving for \(x\) right? Or proving it?

OpenStudy (anonymous):

i think solving for x but not sure.. on the sheet it just says solve.

OpenStudy (loser66):

can I say something? @Mimi_x3 at the first line when you got \[\frac{ 1 }{ \sin^2 x }-1= \csc^2 x -1\]. that's it

OpenStudy (loser66):

ok, sorry for that,. since I continue other's stuff, there is no logic there. now , i start mine

OpenStudy (anonymous):

ok

OpenStudy (mimi_x3):

what does you mean? @Loser66

OpenStudy (mimi_x3):

do*

OpenStudy (loser66):

\[\frac{cotx}{tanx} => \frac{cosx}{sinx} * \frac{1}{tanx} => \frac{cosx}{sinx}* \frac{cosx}{sinx} => \frac{cos^2x}{sin^2x} = \frac{1-sin^2x}{sin^2x} \] = \[\frac{ 1 }{ \sin^2x }-\frac{ \sin^2x }{ \sin^2x }= \csc^2x -1\]

OpenStudy (phi):

mimi your very first line is correct. that last step is replace 1/sin^2 with csc^2 to get csc^2 -1

OpenStudy (mimi_x3):

ohh yeah..

OpenStudy (phi):

I don't think you "solve" this. You prove the identity is true.

OpenStudy (mimi_x3):

but the OP said "solve"

OpenStudy (phi):

it is an identity

OpenStudy (loser66):

how can it be "solve"? since you prove 2 sides are equal. subtract to get 0 , solve for what?

OpenStudy (anonymous):

yeah the problem stated solve but im gussesing from teachers lectures it identity because we never went over truly solving just identities so far

OpenStudy (anonymous):

only been a few problems I have had that the outcome is either 1 or another form of identity